OpenStudy (anonymous):

Which statement is true about the difference 2 radical 7 - radical of 28? It is rational and equal to -2. It is rational and equal to 0. It is irrational and equal to -2 radical 7. It is irrational and equal to radical 7.

OpenStudy (anonymous):

It is rational and equal to 0.

OpenStudy (anonymous):

\[ 2 \sqrt 7 - \sqrt{28}=2 \sqrt 7 -2 \sqrt{7}=0 \]

OpenStudy (anonymous):

Thank you @eliassaab

OpenStudy (anonymous):

\[ 28 = 4 (7)\\ \sqrt{28} =\sqrt 4 \sqrt 7=2 \sqrt 7 \]

OpenStudy (anonymous):

YW